OpenStudy (anonymous):

need a flower garden with a premiter of 16ft and area of 12ft

OpenStudy (anonymous):

2x+2y=16 x*y=12 y=12/x substitute 2x+24/x=16 multiply by x 2x^2+24=16x 2x^2-16x+24=0 (2x-12)(x-2)=0 x=2, 6 if x=2 y=6 if x=6, y=2 so the sides are 6 and 2

OpenStudy (anonymous):

I would assume the garden has a shape of a rectangle, then \(2l+2w=16\) and \(lw=12\). Now solve for \(l\) and \(w\) to find the dimensions of the garden.

OpenStudy (anonymous):

perimeter of 16ft by area of 16ft